OpenStudy (anonymous):

What is the volume of a triangular prism with a height of 3 centimeters and a base that is an equilateral triangle with sides of 4 centimeters? Answer

OpenStudy (anonymous):

wot u think ?

OpenStudy (anonymous):

These were the choices: 24cm cubed 12*sqrt*3cm cubed 48 cm cubed 8*sqrt*3 cm cubed

OpenStudy (campbell_st):

The formula is \[V=\frac{1}{3} Base Area \times height\] For the triangular base  find the height of the equilateral triangle by using Pythagoras' thorem h^2 = 4^2 - 2^2 \[h = \sqrt{12}\] \[h = 2\sqrt{3}\] BAse Area the triangle is \[A = 1/2\times 4 \times2 \sqrt{3} = 4\sqrt{3}\] Volume is \[V = \frac{1}{3} \times 4\sqrt{3} \times3\]

OpenStudy (campbell_st):

ooop... thought it said pyramid but here is the prism information Volume = cross section area x length \[V = 4\sqrt{3} \times 3 = 12\sqrt{3}\]

OpenStudy (anonymous):

thank you